# Scalable Quantum Simulations of Scattering in Scalar Field Theory on 120 Qubits

### v1

@@ VI Summary and outlook
Quantum simulations hold the potential of reliably uncovering the dynamics of complex inelastic collision events, such as those that take place in extremely dense and hot environments. As a first step toward such computations, this work simulates elastic scattering in one-dimensional lattice scalar field theory. Scalable variational quantum algorithms are developed for both state preparation and time evolution, paving the path for future simulations at larger scales. The simulation begins by using a translationally-invariant SC-ADAPT-VQE ansatz to prepare the vacuum of the theory. Wavepackets representing particles are then created on top of the vacuum using a brickwall circuit, and a translationally-invariant brickwall ansatz is used to implement time evolution of the system. The scalability of all circuits in these simulations is a result of the physical properties of states in the system, namely their translational invariance, locality, and the presence of a mass gap. The circuits are determined classically on modest system sizes, and extrapolated to arbitrarily large system sizes to be run on quantum devices. These methods are used to simulate the collision of two wavepackets using 120 qubits of IBM’s 156 qubit superconducting quantum computer ibm_fez. The use of error mitigation and suppression techniques is crucial for the extraction of results that qualitatively agree with classical MPS simulations. Signatures of scattering and effects of the interaction strength are clearly identified in the vacuum-subtracted expectation value of the field determined from data from the quantum computer.
@@
Overall, the utility of variational quantum algorithms in systematically compressing circuits is clear; the specialization of these methods to quantum simulations enables the use of physics to provide further performance improvements. These newly-developed scalable variational methods, together with existing techniques and new modifications to error mitigation strategies, enable the first simulation of a relativistic particle scattering event on a quantum computer. Furthermore, these results establish the first demonstration of the extraction of qualitatively correct observables from quantum devices using variational brickwall circuits at scale. This work establishes variational algorithms as tools that are useful for the compression of all parts of a quantum simulation. In particular, it is demonstrated that the problem of barren plateaus that plagues many variational quantum computations is reduced through knowledge of the system being simulated (e.g., its symmetries and mass gap). Efficient variational circuits are determined by using physics to remove redundancies in the circuit parameterization. The development of variational circuit compression methods that are both scalable and amenable to error mitigation is useful for minimizing noise in the quantum device through reduced two-qubit gate depth. These methods and results constitute an important stepping stone toward the quantum simulation of inelastic scattering processes at high energies, where known classical methods are expected to fail.
@@
Highly-inelastic scattering events are believed to generate a large amount of entanglement, and as a result are prime candidates for demonstrations of quantum utility. Toward the simulation of inelastic collisions, several improvements to the current algorithm will be necessary. Particles with energy higher than the two-particle threshold will need to be created, requiring both a finer lattice spacing $a$, and more tightly peaked wavepackets in $k$-space. In turn, both of these will require wavepackets that occupy a larger number of qubits. The methods presented will need to be extended to the scalable creation of wavepackets of increasing sizes and their time evolution. The presence of a large number of lattice sites occupied by the vacuum will continue to provide the resolution in lattice momentum that will be necessary to observe inelastic effects. More qubits per site (i.e., increasing $n_{q}$) will be needed to achieve higher precision in these calculations. All of these improvements will require circuits with greater two-qubit gate depths. The scalability of the presented methods can also be applied by extrapolating the variational parameters to continuum physics, through taking the continuum limit in the process of the variational optimization. As the capabilities of state of the art quantum computers to realize large-scale system sizes improve, more sophisticated techniques for particle detection will be necessary. Further in the future, the extension of these methods to include more complex interactions, such as those described by non-Abelian gauge theories, will be required for the simulation of realistic particle collisions.
@@ Acknowledgments
The author thanks Roland Farrell, Marc Illa, Zhiyao Li, Henry Froland, Anthony Ciavarella, and Martin Savage for helpful discussions and insightful comments. This work was supported in part by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, InQubator for Quantum Simulation (IQuS) [163] under Award Number DOE (NP) Award DE-SC0020970 via the program on Quantum Horizons: QIS Research and Innovation for Nuclear Science. This work was also supported, in part, through the Department of Physics [164] and the College of Arts and Sciences [165] at the University of Washington. This work has made extensive use of Wolfram Mathematica [166], python [167], julia [168], jupyter [169] notebooks in the conda [170] environment, IBM’s quantum programming environment qiskit [151], and iTensor [171, 172] in this work. This work was enabled, in part, by the use of advanced computational, storage and networking infrastructure provided by the Hyak supercomputer system at the University of Washington [173]. This research was done using services provided by the OSG Consortium [174, 175, 176, 177], which is supported by the National Science Foundation awards #2030508 and #1836650. The author acknowledges the use of IBM Quantum Credits for this work. The views expressed are those of the author, and do not reflect the official policy or position of IBM or the IBM Quantum team.
Appendix A Comparison of exact and variational time evolution
@@
Figure 12: The vacuum-subtracted $\langle\phi^{2}_{j}\rangle$ for different times $t=0-9$ computed exactly using matrix exponentiation (dashed lines) compared to its approximation using SVC (solid lines) for (a) the free $(\lambda=0)$ theory and (b) the interacting $(\lambda=2)$ theory. An $L=12$ (24 qubits) system is used.
@@
In this appendix, noiseless simulations of an $L=12$ system are used to compare exact time evolution in the digitized theory with $n_{q}=2$ to time evolution approximated by SVC. Exact calculations are done using matrix exponentiation applied to the $t=0$ state prepared via SVC methods. Figure 12 shows the comparison for $\lambda=0$ and $\lambda=2$. The SVC variational results can be seen to qualitatively track the exact time evolution, but degrade with time. The quality of the $\lambda=0$ variationally-determined observables is seen to be consistently worse than those for $\lambda=2$, which is again a consequence of the $\lambda=0$ theory having a longer correlation length. Although the local infidelity spanning the interaction region $I_{10}$ is high (see Table 1), the observables calculated from the SVC states still follow the exact values, suggesting that fidelity is a rough bound on the quality of prepared variational states. Since quantum devices compute observables, even states that have high infidelity may be used to approximately simulate the desired process.
@@ Appendix B Optimization
The time evolution circuit optimization is done in a sequential manner over the simulation times. Because of the Trotter step-like nature of the ansatz, results from optimizations of previous steps are used as initial guesses for future optimizations. For example, since $t=4$ uses two variational steps, and $t=2$ uses one step, the angles implementing the $t=2$ evolution are used as initial guesses for the $t=4$ optimization. Table 3 specifies how the initial guesses utilize parameters that have been optimized for past simulation times. The BFGS gradient-based optimization method [178] is used, with a tolerance of $10^{-5}$.
$t$ of variational steps Initial guess $t$ of variational steps Initial guess $t$ of variational steps Initial guess
@@
Appendix C Error mitigation
A suite of error mitigation techniques is used in this work, enabling extraction of meaningful results from noisy data from quantum devices. The central ingredient of the error mitigation strategy used in this work is ODR, which is a local version of decoherence renormalization [156, 185, 186]. In this method, known noise-free results of circuits that are easily simulated classically are compared to results of the same circuits from the quantum device. This ratio is then used in a spatially-local (operator-wise) fashion to measure the noise that has impacted each local expectation value. ODR assumes a depolarizing noise model. The standard method of PT is used to convert the coherent noise in each circuit to Pauli noise, $\rho\rightarrow\sum_{i}p_{i}P_{i}\rho P_{i}$, where $\rho$ is a reduced state used to measure a local observable. As is standard for similar error mitigation strategies, depolarizing noise is assumed, which can be seen as an special case of Pauli noise where all Pauli error terms are equally likely, $\rho\rightarrow(1-p)\rho+p\mathbbm{1}$. The error mitigation scheme works by estimating the parameter $p_{j}$ for this channel for each local observable $O_{j}$.

### v2

@@ VI Summary and outlook
Quantum simulations hold the potential of reliably uncovering the dynamics of complex inelastic collision events, such as those that take place in extremely dense and hot environments. As a first step toward such computations, this work simulates elastic scattering in one-dimensional lattice scalar field theory. Scalable variational quantum algorithms are developed for both state preparation and time evolution, paving the path for future simulations at larger scales. The simulation begins by using a translationally-invariant SC-ADAPT-VQE ansatz to prepare the vacuum of the theory. Wavepackets representing particles are then created on top of the vacuum using a brickwall circuit, and a translationally-invariant brickwall ansatz is used to implement time evolution of the system. The scalability of all circuits in these simulations is a result of the physical properties of states in the system, namely their translational invariance, locality, and the presence of a mass gap. The circuits are determined classically on modest system sizes, and extrapolated to arbitrarily-large system sizes to be run on quantum devices. These methods are used to simulate the collision of two wavepackets using 120 qubits of IBM’s 156 qubit superconducting quantum computer ibm_fez. The use of error mitigation and suppression techniques is crucial for the extraction of results that qualitatively agree with classical MPS simulations. Signatures of scattering and effects of the interaction strength are clearly identified in the vacuum-subtracted expectation value of the field determined from data from the quantum computer.
@@
Overall, the utility of variational quantum algorithms in systematically compressing circuits is clear; the specialization of these methods to quantum simulations enables the use of physics to provide further performance improvements. These newly-developed scalable variational methods, together with existing techniques and new modifications to error mitigation strategies, enable the first simulation of wavepacket scattering in an interacting quantum field theory on a quantum computer. The resulting gate counts and resource requirements are many orders of magnitude below upper bounds for fault-tolerant algorithms. Furthermore, these results are the first demonstration of the extraction of qualitatively correct observables from quantum devices using variational brickwall circuits at scale. This work establishes variational algorithms as tools that are useful for the compression of all parts of a quantum simulation. In particular, it is demonstrated that the problem of barren plateaus that plagues many variational quantum computations is reduced through knowledge of the system being simulated (e.g., its symmetries and mass gap). Efficient variational circuits are determined by using physics to remove redundancies in the circuit parameterization. The development of variational circuit compression methods that are both scalable and amenable to error mitigation is useful for minimizing noise in the quantum device through reduced two-qubit gate depth. These methods and results constitute an important stepping stone toward the quantum simulation of inelastic scattering processes at high energies, where known classical methods are expected to fail.
@@
Highly-inelastic scattering events are believed to generate a large amount of entanglement, and as a result are prime candidates for demonstrations of quantum utility. Toward the simulation of inelastic collisions, several improvements to the current algorithm will be necessary. Initial states with energy higher than the two-particle threshold will need to be created, requiring more tightly peaked wavepackets in $k$-space. In turn, wavepackets that occupy a larger number of qubits will be required. The methods presented will need to be extended to the scalable creation of wavepackets of increasing sizes and their time evolution. The presence of a large number of lattice sites occupied by the vacuum will continue to provide the resolution in lattice momentum that will be necessary to observe inelastic effects. More qubits per site (i.e., increasing $n_{q}$) will be needed to achieve higher precision in these calculations, as well as to simulate genuine interactions. Similarly, larger system sizes and correlation lengths will be needed to approach continuum physics. All of these changes will require circuits with greater two-qubit gate depths, and improvements to the variational methods will be necessary. As the capabilities of state of the art quantum computers to realize large-scale system sizes improve, more sophisticated techniques for particle detection will be required. Further in the future, the extension of these methods to include more complex interactions, such as those described by non-Abelian gauge theories, will be required for the simulation of realistic particle collisions.
@@ Acknowledgments
The author thanks Roland Farrell, Marc Illa, Zhiyao Li, Henry Froland, Anthony Ciavarella, and Martin Savage for helpful discussions and insightful comments. This work was supported in part by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, InQubator for Quantum Simulation (IQuS) [171] under Award Number DOE (NP) Award DE-SC0020970 via the program on Quantum Horizons: QIS Research and Innovation for Nuclear Science. This work was also supported, in part, through the Department of Physics [172] and the College of Arts and Sciences [173] at the University of Washington. This work has made extensive use of Wolfram Mathematica [174], python [175], julia [176], jupyter [177] notebooks in the conda [178] environment, IBM’s quantum programming environment qiskit [158], iTensor [179, 180], and quimb [181] in this work. This work was enabled, in part, by the use of advanced computational, storage and networking infrastructure provided by the Hyak supercomputer system at the University of Washington [182]. This research was done using services provided by the OSG Consortium [183, 184, 185, 186], which is supported by the National Science Foundation awards #2030508 and #1836650. The author acknowledges the use of IBM Quantum Credits for this work. The views expressed are those of the author, and do not reflect the official policy or position of IBM or the IBM Quantum team.
Appendix A Digitization effects
The Nyquist-Shannon Sampling Theorem [115, 116, 117, 118] guarantees that with a proper choice of $\phi_{\text{max}}$, the digitized field theory represents the continuous lattice field theory with errors that are exponentially suppressed with an increasing number of wavefunction sample points. In particular, the eigenfunctions of the continuous theory can be reconstructed from the digitized eigenfunctions up to errors that scale as $\epsilon\sim 2^{-2^{n_{q}}}$. The single-site wavefunction $\psi(\phi)$ can be approximately reconstructed from $n$ samples at points $\phi_{n}$ using the Whittaker–Shannon interpolation formula [115]:
$\psi(\phi)$ $=\sum_{n=0}^{2^{n_{q}}-1}\psi(\phi_{n})\,\text{sinc}\left(\frac{\phi+\phi_{\text{max}}-n\delta_{\phi}\phi_{\text{max}}}{\delta_{\phi}\phi_{\text{max}}}\right).$ (27)
Its multidimensional extensions can be used to reconstruct multi-site wavefunctions.
A poorly chosen $\phi_{\text{max}}$ can introduce unwanted interactions due to the digitization and break the double-exponential convergence guaranteed by the Nyquist-Shannon Sampling Theorem. The optimal choice of $\phi_{\text{max}}$ for a given theory depends on $n_{q}$, as well as on the parameters $m$ and $\lambda$, since these control the strength of the potential. As a result, states with higher $m$ and $\lambda$ are more localized in $\phi$-space and less localized in $\Pi$-space. Previous approaches determined the optimal $\phi_{\text{max}}$ by minimizing the error in the commutator compared to the continuum value $\phi_{i},\Pi_{j}]-i\delta_{ij}|| [$116, 117, 118, 119, 120], or minimizing the error in the eigenenergies by numerically solving the continuum theory for a small number of lattice sites [85]. However, single-site considerations do not capture additions to the potential from neighboring sites, and numerical solutions become intractable for large systems due to the exponential resources required in the discretization of the system. In this work, $\phi_{\text{max}}$ is approximately chosen by including the single-site contributions from $H_{\text{kin}}$, so that the only term that is omitted from the Hamiltonian of Eq. (1) is $\sum_{j=0}^{L-1}\phi_{j+1}\phi_{j}$. Similar to the previous methods, this is an approximation to the optimal $\phi_{\text{max}}$ for large system sizes, but it is closer to the true value. [^21] In principle, the optimal $\phi_{\text{max}}$ can be determined following an approach similar to that of Sec. III, by finding $\phi_{\text{max}}$ as a function of the system size $L$ for small $L$, and extrapolating to $L$ of choice. Following these considerations, the optimal $\phi_{\text{max}}$ is determined by using Eq. (27) to maximize the overlap between the low-energy eigenstates in the digitized theory and the continuum theory solved numerically. With this method, the best choice for $\phi_{\text{max}}$ is 1.5 for $n_{q}=2$, $m=1/2$, $\lambda=0$, and 1.43 for $n_{q}=2$, $m=1/2$, $\lambda=2$. However, for $\lambda=2$ the difference between $\phi_{\text{max}}=1.5$ and the ideal value is minimal in terms of the effect on the representation of the continuum, and as a result $\phi_{\text{max}}=1.5$ is used for both the free and interacting theories throughout this work. In this case, the errors due to the suboptimal choice of $\phi_{\text{max}}$ are small compared to digitization errors stemming from a small $n_{q}$.
Two qubits per lattice site is the coarsest nontrivial digitization. As a result, the digitization effects are significant and must be considered. With $n_{q}=2$ and $\delta_{\phi}$ chosen as in Eq. (13), the $\phi^{2}_{j}$ and $\phi^{4}_{j}$ operators may be written as
$\phi^{2}_{j}=\frac{\phi_{\text{max}}^{2}}{9}\left(5+4Z_{2j}Z_{2j+1}\right),$ (28)
$\phi^{4}_{j}=\frac{\phi_{\text{max}}^{4}}{81}\left(41+40Z_{2j}Z_{2j+1}\right).$ (29)
From these expressions, it is seen that the non-identity portions of $H_{\phi}$ and $H_{\text{int}}$ in the Hamiltonian of Eq. (1) are proportional to each other. This implies that simulating a theory with $\lambda\neq 0$ is equivalent to simulating one with a larger mass,
$m^{\prime}=\sqrt{m^{2}+\frac{\lambda}{4!}\frac{20}{9}\phi_{\text{max}}^{2}}.$ (30)
As a result of this, the interactions present in the simulations in this work are due to digitization effects. In principle, these interactions may be mapped to an expansion of the form $\sum_{n}c_{2n}\phi^{2n}$, but determining the coefficients $c_{2n}$ is difficult in practice. These interactions are examined in Fig. 12, where the time evolution of a single wavepacket is compared to that of a two-wavepacket state for both values of the coupling. By considering the single-wavepacket evolution, which is truly noninteracting, it can be seen that there is a time delay due to the interaction of two wavepackets for both couplings, with the time delay for the $\lambda=2$ theory being larger.
Figure 12: Comparison of the time evolution of the single- and two-wavepacket states ($|\psi_{\text{1wp}}\rangle$ and $|\psi_{\text{2wp}}\rangle$ respectively), for (a) the free $(\lambda=0)$ and (b) the interacting $(\lambda=2)$ theories. The center of the left wavepacket is tracked for the single-wavepacket state with the dotted vertical line, and for the two-wavepacket state with the dash-dotted vertical line. The centers are determined by fitting a Gaussian around the wavepacket region, and by taking the maximum value when the wavepackets are interacting. A system of $L=20$ spatial sites with $n_{q}=2,m=1/2$ is used. The simulations are done using a MPS circuit simulator with size $1/10$ Trotter steps and a maximum bond dimension of 100.
Figure 13: The center $j_{*}$ of the left wavepacket propagating in time with (a) $n_{q}=2$ and (b) $n_{q}=3$ digitization. The free and interacting theories are shown in blue and yellow, respectively. The single-wavepacket states $|\psi_{\text{1wp}}\rangle$ are marked with dashed lines, while the two-wavepacket states $|\psi_{\text{2wp}}\rangle$ are marked with solid lines. A system of $L=20$ spatial sites with $m=1/2$ is used. For $n_{q}=2$, $\phi_{\text{max}}=1.5$ is used for both $\lambda=0$ and $\lambda=2$ as in the simulations run on quantum computers in this work, while for $n_{q}=3$, the optimal $\phi_{\text{max}}$ values of 2.52 and 2.43 are used for the free and interacting theories, respectively. The simulations are done using a MPS circuit simulator with size $t=1/10$ Trotter steps and maximum bond dimension of 100.
The centers $j_{*}$ of the wavepackets for all four cases are plotted in Fig. 13a. The dashed lines show the single-wavepacket time evolution. The slight difference in group velocity $v_{k}$, as expected from Fig. 5, is seen in the slopes of the $|\psi_{\text{1wp}}\rangle$ lines. By comparing the single-wavepacket and the two-wavepacket lines, the relative time delay between the $\lambda=0$ and $\lambda=2$ theories as a result of differences in $v_{k}$ is eliminated. The collision takes place between times t=36t=3-6 indicated by the plateau, and the time delay is seen in the $\lambda=2$ plateau extending later in time than the $\lambda=0$ one. After an initial settling time, the “asymptotic” time delay is also seen in the difference in the center points at a given time. The difference of $j_{*}$ between the $|\psi_{\text{1wp}}\rangle$ and $|\psi_{\text{2wp}}\rangle$ is larger for $\lambda=2$, indicating that the time delay due to interactions is greater in this theory.
To complete the discussion of digitization effects and interactions caused by them, simulations of the same process with $n_{q}=3$ are shown in Fig. 13b. These simulations are expected to more accurately capture the continuum physics because of the added precision from doubling the number of points where the continuum wavefunctions are sampled. In addition, the non-identity parts of $H_{\phi}$ and $H_{\text{int}}$ are not proportional with $n_{q}=3$, making the $\phi^{4}$ operator a genuine interaction term not present in the free Hamiltonian. Using the methods described above, the optimal $\phi_{\text{max}}$ is determined to be 2.52 for $\lambda=0$, and 2.43 for $\lambda=2$. All other parameters of the simulations are the same as for $n_{q}=2$. In the $\lambda=0$ simulations for $n_{q}=3$ there is no time delay, and a very small time delay for $\lambda=2$, which shows the strength of the digitization effects for $n_{q}=2$.
Appendix B Details on wavepacket creation
Section IV.2 describes wavepacket preparation by learning an approximate shallow-depth variational circuit. This circuit is optimized to maximize the overlap with $|\psi_{\text{targ}}\rangle$ being the wavepacket produced by the non-digitized analytical wavefunction (Eq. (10)) for $\lambda=0$, and after adiabatically turning on the interactions (Eq. (12)) for $\lambda=2$. The wavepackets represented by $|\psi_{\text{targ}}\rangle$ above are compared to those determined from ED by block-diagonalizing the Hamiltonian into momentum sectors in Table 3. The local infidelity over the three sites spanning the wavepacket, $I_{3}$, is seen to decrease with increasing $n_{q}$ indicating that the digitized wavefunction approximates the continuous-field case with increasing quality. [^22] The wavepackets produced from ED can also be used directly as $|\psi_{\text{targ}}\rangle$, eliminating the need for the continuous-field approximation (Eq. (10)) and adiabatic turn-on (Eq. (12)). This approach would encounter the same classical limitations in terms of scalability and is conceptually simpler. However, the direct ED method was not used in determining the parameters for circuits run on quantum hardware in this work.
$I_{3}\,\,(n_{q}=2)$ $I_{3}\,\,(n_{q}=3)$
$\lambda=0$ 0.03922 0.01471
$\lambda=2$ 0.03973 0.01537
Table 3: Comparison of wavepackets produced from the continuous-field approximation (Eq. (10)) with adiabatic turn-on using Eq. (12) for $\lambda=2$ to those determined from ED, as a function of the number of qubits used in the digitization $n_{q}$. The free and interacting theories are shown in the center and bottom rows, respectively. The wavepackets are created on an $L=6$ system for both methods. The local infidelity over the three sites spanning the wavepacket, $I_{3}$ is used. For $n_{q}=\{2,3\}$, the values of $\phi_{\text{max}}=\{1.5,3.1\}$ are used as determined in Ref. [85].
The adiabatic turn-on is implemented as follows. The time evolution is implemented using Eq. (12) for a total time of $t_{\text{ad}}=100$, broken into $N_{\text{ad}}=10=O(\sqrt{\lambda t_{\text{ad}}})$ steps (as determined in Refs. [83, 84]). Each adiabatic step in Eq. (12) is Trotterized into 1000 “forward” and 1000 “backward” steps. The forward steps linearly increase the value of $\lambda$ once every 10 steps, while the backward steps evolve with fixed $\lambda$. As explained in Sec. II, the backward evolution steps serve to undo the unwanted time evolution to individual eigenstates composing the wavepacket.
The profile of $|\psi_{\text{targ}}\rangle$, and thus of the produced wavepackets is specified to be a Gaussian in $k$-space, as in Eqs. (9) and (10). This is slightly different from the approach JLP take in Refs. [83, 84], where the form of the state is specified in position space. By specifying a Gaussian in $k$-space, the method used in this work also specifies a Gaussian in position space, where its size is then truncated to optimize device resources. The two approaches are equivalent, since JLP also truncate the spatial extent of the wavepackets.
Appendix C Comparison of exact and variational time evolution
@@
Figure 14: The vacuum-subtracted $\langle\phi^{2}_{j}\rangle$ for different times t=09t=0-9 computed using a MPS circuit simulator for Trotterized time evolution (dashed lines) compared to its approximation using SVC (solid lines) for (a) the free $(\lambda=0)$ theory and (b) the interacting $(\lambda=2)$ theory. An $L=20$ (40 qubits) system is used with maximum bond dimension 100. The exact results use Trotterization with size $1/10$ Trotter steps.
@@
In this appendix, noiseless MPS circuit simulations of an $L=20$ system are used to compare exact time evolution in the digitized theory with $n_{q}=2$ to time evolution approximated by SVC. Exact calculations are done using matrix exponentiation applied to the $t=0$ $|\psi_{\text{2wp}}\rangle$ state prepared via SVC methods. Figure 14 compares $\lambda=0$ and $\lambda=2$. The SVC variational results can be seen to qualitatively track the exact time evolution, but degrade with time. The quality of the $\lambda=0$ variationally-determined observables is seen to be consistently worse than those for $\lambda=2$, which is again a consequence of the $\lambda=0$ theory having a longer correlation length. Although the local infidelity spanning the interaction region $I_{10}$ is high (see Table 1), the observables calculated from the SVC states still follow the exact values, suggesting that fidelity is a rough bound on the quality of prepared variational states. Since quantum devices compute observables, even states that have high infidelity may be used to approximately simulate the desired process.
@@ Appendix D Optimization
The time evolution circuit optimization is done in a sequential manner over the simulation times. Because of the Trotter step-like nature of the ansatz, results from optimizations of previous steps are used as initial guesses for future optimizations. For example, since $t=4$ uses two variational steps, and $t=2$ uses one step, the angles implementing the $t=2$ evolution are used as initial guesses for the $t=4$ optimization. Table 4 specifies how the initial guesses utilize parameters that have been optimized for past simulation times. The Broyden–Fletcher–Goldfarb–Shanno (BFGS) gradient-based optimization method [187] is used, with a tolerance of $10^{-5}$.
$t$ # of variational steps Initial guess $t$ # of variational steps Initial guess $t$ # of variational steps Initial guess
@@
Appendix E Error mitigation
A suite of error mitigation techniques is used in this work, enabling extraction of meaningful results from noisy data from quantum devices. The central ingredient of the error mitigation strategy used in this work is ODR, which is a local version of decoherence renormalization [163, 194, 195]. In this method, known noise-free results of circuits that are easily simulated classically are compared to results of the same circuits from the quantum device. This ratio is then used in a spatially-local (operator-wise) fashion to measure the noise that has impacted each local expectation value. ODR assumes a Pauli noise model. The standard method of PT is used to convert the coherent noise in each circuit to Pauli noise, $\rho\rightarrow\sum_{i}p_{i}P_{i}\rho P_{i}$, where $\rho$ is a reduced state used to measure a local observable. The expectation value of an observable $O$ under this noise channel is given by $\langle O\rangle=\sum_{i}p_{i}\text{Tr}(P_{i}\rho P_{i}O)$. For Pauli observables, $P_{i}OP_{i}=\pm O$, and this channel rescales the measured values $\langle O\rangle_{\text{meas}}$ compared to their noise-free values $\langle O\rangle_{\text{true}}$, $\langle O\rangle_{\text{meas}}=p\langle O\rangle_{\text{true}}$. The error mitigation scheme works by estimating the parameter $p_{j}$ for this channel for each local observable $O_{j}$.